The pH of blood plasma is 7.40. The principal buffer system is HCO3?/H2CO3. Calculate the ratio [HCO3?]/[H2CO3] in blood plasma. (Ka of H2CO3, carbonic acid, is 4.5 × 10?7) Enter your answer in the provided box.
Answer #1

The pH of blood constituents of the bicarbonate buffering system, pKa and ratio of Bicarbonate to carbonic acid is releted to the Henderson–Hasselbalch equation as:

pH = pKacarbonic acid + log10 [HCO-3 / H2CO3]

Since pKa = - log10 Ka

Buffer reaction is : H2CO3  \rightarrow? HCO-3 + H+

Steps:

Given pH of blood plasma: 7.40

Ka of H2CO3 : 4.5 X 107

Find :  ratio [HCO3-]/[H2CO3]

1.Put the all values given in H. H. equation then

pH = pKacarbonic acid + log10 [HCO-3 / H2CO3]

7.40 = - log10 (4.5 X 107) + log10 [HCO-3 / H2CO3]

7.40 + log10 (4.5 X 107) = log10 [HCO-3 / H2CO3]

7.40 + log10 (4.5) + 7 log10 10 = log10 [HCO-3 / H2CO3]

7.40 + 0.6532 + 7 ( log10 10 = 1) = log10 [HCO-3 / H2CO3]

15.0532 =  log10 [HCO-3 / H2CO3]

log10 [HCO-3 / H2CO3] = 15.0532

[HCO-3 / H2CO3] = Anti log (15.0532)

[HCO-3 / H2CO3] = 1130316324738697.8

hence [HCO-3 / H2CO3] = 11.30 X 1014